A certain compound was found to contain 54.0 g of carbon and 10.5 grams of hydrogen Its relative molecular mass is 86.0 Find the empirical and the molecular formulas


1
Expert's answer
2021-02-15T05:57:10-0500

n(C)=54/12=4,5mol

n(H)=10.5/1=10.5mol


C:H=4,5:10,5=1:2


CH2

M(CH2)=14g/mol


86/14=6


Answer: C6H14